Depmsey Candy company wants to make a new mixture. How many pounds of nut mix worth $8 per pound should be added to 20 pounds of choclate drips worth $5 per pound to get a mixture work $6 per pound?
Answer by Earlsdon(6294)

Let x = the number of pounds of nut mix at $8 per pound.
x($8)+20($5) = (x+20)($6) Simplify and solve for x.
8x+100 = 6x+120 Subtract 6x from both sides.
2x+100 = 120 Subtract 100 from both sides.
2x = 20 Finally, divide both sides by 2.
x = 10
The Dempsey company will need to add 10 pounds of nut mix at $8 per pound to 20 pounds of chocolate drips at $5 per pound to obtain 30 pounds of mixture at $6 per pound.